# Projects/QSM24/SFVisco/ViscoSF.tex

### v4

@@ II Analysis
II.1 Non-local Dynamics
I have previously given a formally exact formulation of quantum statistical mechanics in classical phase space (Attard 2018, 2021, 2023a). Decoherence due to entanglement with the reservoir or environment is essential to the formulation. In this the configuration picture I rely upon the momentum eigenfunction $\phi_{\bf p}({\bf q})=V^{-N/2}e^{-{\bf p}\cdot{\bf q}/\mathrm{i}\hbar}$ (Merzbacher 1970, Messiah 1961). Here for $N$ particles the momentum configuration is ${\bf p}=\{{\bf p}_{1},{\bf p}_{2},\ldots,{\bf p}_{N}\}$, and the position configuration is ${\bf q}=\{{\bf q}_{1},{\bf q}_{2},\ldots,{\bf q}_{N}\}$. The momenta are quantized, with momentum state spacing $\Delta_{p}=2\pi\hbar/L$, with $V=L^{3}$ being the volume of the cubic subsystem, and $\hbar=1.05\times 10^{-34}$ being Planck’s constant divided by $2\pi$. The spacing between momentum states goes to zero in the thermodynamic limit. Note that a point in classical phase space, ${\bm{\Gamma}}\equiv\{{\bf q},{\bf p}\}$, has the interpretation of a specific configuration of bosons at these positions with these (quantized) momenta, and it has associated with it the complex number $\phi_{\bf p}({\bf q})$.
@@ II.1 Non-local Dynamics
Here and throughout the occupancy of the single particle momentum state ${\bf a}$ is $N_{\bf a}=\sum_{j=1}^{N}\delta_{{\bf p}_{j},{\bf a}}$. This is what ultimately drives Bose-Einstein condensation (Attard 2025). The formulation of quantum statistical mechanics relies upon the fact that the subsystem is decoherent due to entanglement with the reservoir or environment.
@@
The reason for neglecting the terms involving permutations of bosons with dissimilar momenta is that these are more or less randomly and uniformly distributed on the unit circle in the complex plane, and so they add up to zero. This is particularly the case when one considers that small changes in the positions may lead to wildly different exponents for any such permutations. The sum that is retained involves only permutations between bosons in the same, or nearly the same, momentum state, in which case the exponent is zero, or close to zero, even for small changes in positions. The number of permutations between bosons in exactly the same momentum states in the double sum is $\sum_{\hat{\mathrm{P}}^{\prime},\hat{\mathrm{P}}^{\prime\prime}}\!^{(\hat{\mathrm{P}}^{\prime}{\bf p}=\hat{\mathrm{P}}^{\prime\prime}{\bf p})}=N!\prod_{\bf a}N_{\bf a}({\bf p})!=N!\chi^{+}_{\bf p}$. Arguably these are the ones that dominate, with the permutations between bosons in neighboring momentum states providing a correction.
@@
$[\hat{\mathrm{I}}+(\tau/\mathrm{i}\hbar)\hat{\cal H}({\bf q})]\phi_{\bf p}({\bf q})=\phi_{\overline{\bf p}^{\prime}}(\overline{\bf q}^{\prime}).$ (2.4)
@@
These are Hamilton’s classical equations of motion. The second says that a boson’s momentum evolves according to the classical force acting on it.
Focussing on the occupied momentum state ${\bf a}$, consecutive permutations of the $N_{\bf a}$ bosons initially in it has over a time $\tau\gtrsim\tau_{\mathrm{mix}}$ the effect of sharing equally the classical forces acting on them. On average they therefore evolve classically together, each under the influence of identical non-local forces. The dynamics of a momentum state are those of a rigid body, except that the bosons respond to the equally shared non-local forces on an individual stochastic basis (see the next section).
In detail, for short time intervals, $\tau<\tau_{\mathrm{mix}}$, the subsystem can exist in a superposition of permutations of nearby momentum states as each evolves according to local forces. Eventually, however, for $\tau>\tau_{\mathrm{mix}}$ this has to be winnowed down to a small window about a single pure state. This state must satisfy $\hat{\mathrm{P}}\overline{\bf p}^{\prime}(\tau|{\bf q},{\bf p})=\overline{\bf p}^{\prime}(\tau|{\bf q},{\bf p})$, for any permutation $\hat{\mathrm{P}}$ such that $\hat{\mathrm{P}}{\bf p}={\bf p}$. This is the state predicted by the non-local force per boson in the initial momentum state. In the random phase analysis of the Born probability above, amongst all the superposed states, the one with the bosons all in the same new momentum state is the one with the greatest weight.
Figure 1: Two bosons moving near repulsive walls. The bosons are initially in the same momentum state (left). After an intermediate transposition (middle), the final configuration with equal momenta (lower right) has double weight due to its identical transpose. The final configuration with different momenta (upper right) partly cancels with its transpose.
To put it another way, the original momentum configuration with $N_{\bf a}$ bosons in the momentum state ${\bf a}$ evolves along initially $N_{\bf a}!$ trajectories according to the classical local forces. These trajectories, or the configurations on them, are superposed. Each configuration in the superposition then bifurcates into $N_{\bf a}!$ further trajectories. Thus the number of superposed trajectories grows exponentially. These trajectories diverge from each other in momentum space, and the most divergent are most quickly culled by the mechanism of decoherence. The trajectories that come to dominate are those that are closest together, which eventually means the one with all the $N_{\bf a}$ bosons in the same momentum state (cf. Fig. 1). This is the trajectory that results from all the bosons sharing equally over the time interval the local forces acting on each of them individually.
Of course from the classical perspective, in the thermodynamic limit the spacing between momentum states becomes infinitesimal, and there is no measurable distinction between the subsystem being in a momentum configuration and being in a superposition of neighboring momentum configurations.
The shared non-local force per boson in momentum state ${\bf a}$ is (Attard 2025)
${\bf F}_{\bf a}=\frac{1}{N_{\bf a}}\sum_{j\in{\bf a}}{\bf f}_{j}.$ (2.6)
Here the local classical force on boson $j$ is ${\bf f}_{j}=-\nabla_{q,j}U({\bf q})$, where $U({\bf q})$ is the potential energy. The permutation of the momenta of the bosons in a momentum state effectively averages the local forces and applies it non-locally across the momentum state.
The occupation entropy, $S^{\mathrm{occ}}({\bf p})=k_{\mathrm{B}}\ln\chi_{\bf p}^{+}$, acts like an additional force that seeks to keep bosons together in the same momentum state. The shared non-local force preserves the occupation entropy on average.
@@
It should be emphasized that there are two sources of decoherence: there is the internal decoherence that was discussed above on the basis of the Born probability, and there is the eternal decoherence due to the entanglement of the open subsystem with the environment or reservoir. Both mechanisms dictate that the subsystem must be in a pure state where each boson has a position and a momentum. Allowed permutations are between bosons in the same momentum state, and, for limited time periods, also between bosons in neighboring momentum states. Permutations between bosons in different momentum states, which would correspond to the superposition of different momentum configurations, are suppressed, with the suppression increasing with increasing distance between the permuted states.
II.2 Adiabatic Transition
@@
The momenta are quantized, with ${\bf p}$ being a $3N$-dimensional vector integer multiple of $\Delta_{p}$.
Consider the transition of a single boson in the momentum state ${\bf a}$, which has $N_{\bf a}$ bosons, to a neighboring momentum state ${\bf a}^{\prime}$, which has $N_{{\bf a}^{\prime}}$ bosons, $\{N_{{\bf a}},N_{{\bf a}^{\prime}}\}\stackrel{{\scriptstyle\tau}}{{\rightarrow}}\{N_{{\bf a}}-1,N_{{\bf a}^{\prime}}+1\}$. The conditional transition probability for the time step gives the transition rate. Bayes theorem for the conditional transition probability with microscopic reversibility is (Attard 2012a)
$\wp(N_{{\bf a}}-1,N_{{\bf a}^{\prime}}+1;t+\tau|N_{{\bf a}},N_{{\bf a}^{\prime}};t)\wp(N_{{\bf a}},N_{{\bf a}^{\prime}};t)$ (2.8)
$=$ $\wp(N_{-{\bf a}},N_{-{\bf a}^{\prime}};t|N_{-{\bf a}}-1,N_{-{\bf a}^{\prime}}+1;t-\tau)$
@@
or
$\frac{\wp(N_{-{\bf a}}-1,N_{-{\bf a}^{\prime}}+1;t-\tau)}{\wp(N_{{\bf a}},N_{{\bf a}^{\prime}};t)}$
$=$ $\frac{\wp(N_{{\bf a}}-1,N_{{\bf a}^{\prime}}+1;t|N_{{\bf a}},N_{{\bf a}^{\prime}};t-\tau)}{\wp(N_{-{\bf a}},N_{-{\bf a}^{\prime}};t+\tau|N_{-{\bf a}}-1,N_{-{\bf a}^{\prime}}+1;t)}.$
If all of the momenta are reversed, ${\bf p}\Rightarrow-{\bf p}$, then $N_{-{\bf a}}=N_{{\bf a}}$ and $N_{-{\bf a}^{\prime}}=N_{{\bf a}^{\prime}}$.
The change in potential energy due to the motion of the single boson under consideration during the time step is
$U({\bf q}(-{\bf a}^{\prime},t-\tau))-U({\bf q})$ (2.10)
$=$ $U({\bf q}+\tau{\bf a}^{\prime}/m)-U({\bf q})$
$=$ $\frac{-\tau}{m}{\bf F}_{{\bf a}}\cdot{\bf a}^{\prime}$
$=$ $\frac{-\tau}{m}{\bf F}_{{\bf a}}\cdot{\bf a}+{\cal O}(\tau\tilde{\Delta}_{p}).$
Notice that this takes boson $j\in{\bf a}$ to experience the shared non-local force per boson, ${\bf F}_{{\bf a}}=N_{\bf a}^{-1}\sum_{k\in{\bf a}}{\bf f}_{k}({\bf q})$, rather than ${\bf f}_{j}$, the local classical force. The total change in potential energy due to bosons in the momentum state ${\bf a}$ obtained with the shared non-local force is the same as would have been obtained summing over the change due to the individual local classical forces.
With this, in the occupancy picture (Attard 2021, 2023a, 2025, Pathria 1972), the left hand side of Bayes’ theorem is
$\frac{\wp(N_{-{\bf a}}-1,N_{-{\bf a}^{\prime}}+1;t-\tau)}{\wp(N_{{\bf a}},N_{{\bf a}^{\prime}};t)}$ (2.11)
$=$ $\frac{\wp(N_{{\bf a}}-1,N_{{\bf a}^{\prime}}+1;t-\tau)}{\wp(N_{{\bf a}},N_{{\bf a}^{\prime}};t)}$
$=$ $e^{-\beta[a^{\prime 2}-a^{2}]/2m}e^{\beta\tau{\bf F}_{{\bf a}}\cdot{\bf a}/m}$
$\approx$ $1-\frac{\beta}{m}[{\bf a}^{\prime}-{\bf a}]\cdot{\bf a}+\frac{\beta\tau}{m}{\bf F}_{{\bf a}}\cdot{\bf a}$
@@
This is the occupancy picture and there is no occupation entropy. Most commonly the transition is to the nearest neighbor state, $\tilde{\Delta}_{p}=\Delta_{p}$. But as discussed below there may be occasions when it is necessary to jump further afield, in which case $\tilde{\Delta}_{p}$ is some integer multiple of $\Delta_{p}$.
For each boson in an occupied state ${\bf a}$, we choose to attempt three successive transitions to a neighbor in the direction of each component of the force, rather than a single transition to one of the seven neighboring states. Also, the shared non-local force per boson ${\bf F}_{{\bf a}}$ does not change during a time step even though the occupancy of the state may change during a time step as all the bosons in the subsystem sequentially attempt a transition.
For the occupied momentum state ${\bf a}$, the relevant neighbor state in the direction $\alpha\in\{x,y,z\}$ is ${\bf a}^{\prime}_{\alpha}\equiv{\bf a}+\mbox{sign}(\tau F_{{\bf a},\alpha})\tilde{\Delta}_{p}\hat{\bf x}_{\alpha}$. (Most commonly the time step $\tau$ is positive and the neighboring momentum state lies in the direction of the force.) The conditional transition probability for the transition $\{N_{{\bf a}},N_{{\bf a}^{\prime}_{\alpha}}\}\stackrel{{\scriptstyle\tau}}{{\rightarrow}}\{N_{{\bf a}}-1,N_{{\bf a}^{\prime}_{\alpha}}+1\}$ may be split into even and odd terms,
$\wp(N_{{\bf a}}-1,N_{{\bf a}^{\prime}_{\alpha}}+1;t|N_{{\bf a}},N_{{\bf a}^{\prime}_{\alpha}};t-\tau)=\wp^{+}_{\alpha}({\bf a})+\wp^{-}_{\alpha}({\bf a}),$ (2.12)
where $\wp^{\pm}_{\alpha}(-{\bf a})=\pm\wp^{\pm}_{\alpha}({\bf a})$. The probability of not changing state is $\wp(N_{{\bf a}},N_{{\bf a}^{\prime}_{\alpha}};t|N_{{\bf a}},N_{{\bf a}^{\prime}_{\alpha}};t-\tau)=1-\wp^{+}_{\alpha}({\bf a})-\wp^{-}_{\alpha}({\bf a})$. With this and the above result Bayes’ theorem reads
[…]
$=$ $\frac{\wp^{+}_{\alpha}({\bf a})+\wp^{-}_{\alpha}({\bf a})}{\wp^{+}_{\alpha}(-{\bf a}^{\prime}_{\alpha})+\wp^{-}_{\alpha}(-{\bf a}^{\prime}_{\alpha})}$
$\approx$ $\frac{\wp^{+}_{\alpha}({\bf a})}{\wp^{+}_{\alpha}({\bf a}^{\prime}_{\alpha})}\left[1+\frac{\wp^{-}_{\alpha}({\bf a})}{\wp^{+}_{\alpha}({\bf a})}+\frac{\wp^{-}_{\alpha}({\bf a}^{\prime}_{\alpha})}{\wp^{+}_{\alpha}({\bf a}^{\prime}_{\alpha})}\right].$
Now assume that to leading order
$\wp^{+}_{\alpha}({\bf a}^{\prime}_{\alpha})=\wp^{+}_{\alpha}({\bf a})[1+\tilde{\Delta}_{p}^{2}b_{\alpha}],$ (2.14)
and
$\wp^{-}_{\alpha}({\bf a}^{\prime}_{\alpha})=\wp^{-}_{\alpha}({\bf a})=\wp^{+}_{\alpha}({\bf a})[\tau c_{\alpha}+\tilde{\Delta}_{p}d_{\alpha}].$ (2.15)
These give
[…]
$=$ $[1-\tilde{\Delta}_{p}^{2}b_{\alpha}]\left[1+2\tau c_{\alpha}+2\tilde{\Delta}_{p}d_{\alpha}\right]$
$=$ $1+2\tau c_{\alpha}+2\tilde{\Delta}_{p}d_{\alpha},$
with second order terms neglected. Hence
$c_{\alpha}=\frac{\beta}{2m}{F}_{{\bf a},\alpha}a_{\alpha},\mbox{ and }d_{\alpha}=\frac{\beta}{2m}\mbox{sign}(\tau F_{{\bf a},\alpha})a_{\alpha}.$ (2.17)
(The sign of $d_{\alpha}$ is corrected in § D. Fortunately, these odd terms are small compared to unity and the numerical results are little affected.) Both of these are odd in $a_{\alpha}$, as required.
The average change in kinetic energy due to the stochastic transition in the $\alpha$ direction is
@@
$\left\langle\Delta_{\alpha}{\cal K}\right\rangle$ $=$ $\frac{\mbox{sign}(\tau F_{{\bf a},\alpha})\tilde{\Delta}_{p}{a}_{\alpha}}{m}$ (2.18)
[…]
@@
$=$ $\frac{\mbox{sign}(\tau F_{{\bf a},\alpha})\tilde{\Delta}_{p}{a}_{\alpha}}{m}[\wp^{+}_{\alpha}({\bf a})+\wp^{-}_{\alpha}({\bf a})]$

### v5

@@ II Analysis
II.1 Hamiltonian Dynamics
Quantum statistical mechanics may be formulated in classical phase space (Attard 2018, 2021, 2023a). This exact transformation relies upon decoherence due to entanglement with the reservoir or environment. This configuration picture invokes the momentum eigenfunction $\phi_{\bf p}({\bf q})=V^{-N/2}e^{-{\bf p}\cdot{\bf q}/\mathrm{i}\hbar}$ (Merzbacher 1970, Messiah 1961). Here for $N$ particles the momentum configuration is ${\bf p}=\{{\bf p}_{1},{\bf p}_{2},\ldots,{\bf p}_{N}\}$, and the position configuration is ${\bf q}=\{{\bf q}_{1},{\bf q}_{2},\ldots,{\bf q}_{N}\}$. The momenta are quantized, with momentum state spacing $\Delta_{p}=2\pi\hbar/L$, with $V=L^{3}$ being the volume of the cubic subsystem, and $\hbar=1.05\times 10^{-34}$ s being Planck’s constant divided by $2\pi$. The spacing between momentum states goes to zero in the thermodynamic limit. Note that a point in classical phase space, ${\bm{\Gamma}}\equiv\{{\bf q},{\bf p}\}$, has the interpretation of a specific configuration of bosons at these positions with these (quantized) momenta, and it has associated with it the complex number $\phi_{\bf p}({\bf q})$.
@@ II.1 Hamiltonian Dynamics
Here and throughout the occupancy of the single particle momentum state ${\bf a}$ is $N_{\bf a}=\sum_{j=1}^{N}\delta_{{\bf p}_{j},{\bf a}}$. This is what ultimately drives Bose-Einstein condensation (Attard 2025).
@@
The reason for neglecting the terms involving permutations of bosons with dissimilar momenta is that these are more or less randomly and uniformly distributed on the unit circle in the complex plane, and so they add up to zero. This is particularly the case when one considers that small changes in the positions may lead to wildly different exponents for any such permutations. The sum that is retained involves only permutations between bosons in the same, or nearly the same, momentum state, in which case the exponent is zero, or close to zero, even for small changes in positions. The number of permutations between bosons in exactly the same momentum states in the double sum is $\sum_{\hat{\mathrm{P}}^{\prime},\hat{\mathrm{P}}^{\prime\prime}}\!^{(\hat{\mathrm{P}}^{\prime}{\bf p}=\hat{\mathrm{P}}^{\prime\prime}{\bf p})}=N!\prod_{\bf a}N_{\bf a}({\bf p})!=N!\chi^{+}_{\bf p}$. Arguably these are the ones that dominate, particularly on the low side of the $\lambda$-transition, which means that we may take $\phi^{+}_{\bf p}({\bf q})=\phi_{\bf p}({\bf q})$.
@@
$[\hat{\mathrm{I}}+(\tau/\mathrm{i}\hbar)\hat{\cal H}({\bf q})]\phi_{\bf p}({\bf q})=\phi_{{\bf p}^{\prime}}({\bf q}^{\prime}).$ (2.4)
@@
These are Hamilton’s classical equations of motion. The second says that a boson’s momentum evolves according to the classical force acting on it, $\dot{\bf p}_{j}=-\nabla_{q,j}{\cal H}({\bf q},{\bf p})={\bf f}_{j}$. Here it is implicitly assumed that the spacing between momentum states is small enough to allow the continuum approximation.
@@
It should be emphasized that there are two sources of decoherence: there is the internal decoherence that was discussed above on the basis of the Born probability, and there is the external decoherence due to the entanglement of the open subsystem with the environment or reservoir. Both mechanisms dictate that the subsystem must be in a pure state where each boson has a position and a momentum. Allowed permutations are between bosons in the same momentum state. Those permutations which would correspond to the superposition of different momentum configurations are suppressed.
II.2 Newton’s Second Law: The Condensed Version
For a momentum configuration ${\bf p}$ in the condensed regime, the single-particle quantized momentum state ${\bf a}$ is occupied by $N_{\bf a}=\sum_{j=1}^{N}\delta_{{\bf p}_{j},{\bf a}}$ bosons. Let the $n_{A}$ bosons in a specific subset $A\in{\bf a}$ be $A=\{j_{1},j_{2},\ldots,j_{n_{A}}\}$, with ${\bf p}_{j_{k}}={\bf a}$. The momentum eigenfunction for this subset is $\phi_{{\bf p}^{n_{A}}}({\bf q}^{n_{A}})$. Let $\hat{\mathrm{P}}_{\!A}$ be one of the $n_{A}!$ permutations amongst the bosons in the specific subset $A$. Obviously the momentum configuration is unchanged by such a permutation, $\hat{\mathrm{P}}_{\!A}{\bf p}^{n_{A}}={\bf p}^{n_{A}}={\bf a}^{n_{A}}$. The Schrödinger time propagator for this subset yields
$\left[\hat{\mathrm{I}}+\frac{\tau}{\mathrm{i}\hbar}\hat{\cal H}({\bf q}^{n_{A}})\right]\phi_{{\bf p}^{n_{A}}}({\bf q}^{n_{A}})$ $=$ $\phi_{{\bf p}^{\prime n_{A}}}({\bf q}^{\prime n_{A}})$
$=$ $\frac{1}{n_{A}!}\sum_{\hat{\mathrm{P}}_{\!A}}\phi_{\hat{\mathrm{P}}_{\!A}{\bf p}^{\prime n_{A}}}({\bf q}^{\prime n_{A}}),$
since the left hand side is unchanged by the sum over permutations. This suggests that the subset moves rigidly, which is to say ${\bf p}_{j_{k}}^{\prime}={\bf a}^{\prime}_{A}$ for ${j_{k}}\in A$, since this ensures that $\hat{\mathrm{P}}_{\!A}{\bf p}^{\prime n_{A}}={\bf p}^{\prime n_{A}}$.
The classical equations of motion become
${\bf p}^{\prime n_{A}}$ $=$ $\frac{1}{n_{A}!}\sum_{\hat{\mathrm{P}}_{\!A}}\hat{\mathrm{P}}_{\!A}{\bf p}^{\prime n_{A}}$ (2.7)
$=$ ${\bf p}^{n_{A}}+\frac{\tau}{n_{A}!}\sum_{\hat{\mathrm{P}}_{\!A}}\hat{\mathrm{P}}_{\!A}{\bf f}^{n_{A}}$
$=$ ${\bf p}^{n_{A}}+\tau{\bf F}_{A}^{n_{A}}.$
It is obvious that symmetrizing the force by permutation over the subset is equivalent to the average non-local force per boson for the subset, ${\bf F}_{A}\equiv(n_{A}!)^{-1}\sum_{\hat{\mathrm{P}}_{\!A}}\big{\{}\hat{\mathrm{P}}_{\!A}{\bf f}^{n_{A}}\big{\}}_{j_{k}}=n_{A}^{-1}\sum_{j\in A}{\bf f}_{j}$.
There are ${}^{N_{\bf a}}C_{n_{A}}\equiv N_{\bf a}!/(N_{\bf a}-n_{A})!n_{A}!$ ways of allocating labeled bosons from the momentum state ${\bf a}$ with $N_{\bf a}$ bosons to a subset $A$ with $n_{A}$ bosons. Label these $n=1,2,\ldots,^{N_{\bf a}}\!C_{n_{A}}$. Each one of these has an indistinguishable initial configuration, ${\bf a}^{n_{A}}$, but a different average non-local force ${\bf F}_{A_{n}}$ and hence a different evolved configuration ${\bf p}^{\prime n_{A}}_{n}$. The evolved configuration is a superposition of these. But as an open quantum system, it must collapse into a single configuration. The specific configuration occurs with probability $1/\,^{N_{\bf a}}C_{n_{A}}$, which means that the quantum equation of motion for the evolution of the specific subset of $n_{A}$ bosons in the momentum state ${\bf a}$ is
${\bf p}^{\prime n_{A}}={\bf p}^{n_{A}}+\tau\frac{(N_{\bf a}-n_{A})!n_{A}!}{N_{\bf a}!}{\bf F}_{A}^{n_{A}},$ (2.8)
where the $A$ refers to a specific labeled subset of the ${}^{N_{\bf a}}C_{n_{A}}$ possible subsets.
We can confirm that this is correct and gain some insight into the physical process by looking at energy conservation. If the momentum state is divided into the subset $A$ with $n_{A}$ bosons, and $A^{\prime}$ with $n_{A^{\prime}}=N_{\bf a}-n_{A}$ bosons, then the respective forces per boson are ${\bf F}_{A}$ and ${\bf F}_{A^{\prime}}$. The total force, $n_{A}{\bf F}_{A_{n}}+n_{A^{\prime}}{\bf F}_{A^{\prime}_{n}}=\sum_{j\in a}{\bf f}_{j}$, is the same for each of the ${}^{N_{\bf a}}C_{n_{A}}$ possible $A_{n}$. The change in total potential energy over the time step for the bosons in ${\bf a}$ is $\Delta U=-\tau\sum_{j\in a}{\bf f}_{j}\cdot{\bf a}/m$ because there is just one position configuration. With the above condensed equation of motion the change in total kinetic energy for the superposed momentum states is
$\Delta{\cal K}$ $=$ $\frac{1}{m}\sum_{n=1}^{{}^{N_{\bf a}}\!C_{n_{A}}}\Big{\{}{\bf a}^{n_{A}}\cdot[{\bf p}^{\prime n_{A}}_{n}-{\bf a}^{n_{A}}]$ (2.9)
[…]
$=$ $\frac{\tau}{m}\frac{n_{A^{\prime}}!n_{A}!}{N_{\bf a}!}\sum_{n=1}^{{}^{N_{\bf a}}\!C_{n_{A}}}\Big{\{}n_{A}{\bf a}\cdot{\bf F}_{A_{n}}+n_{A^{\prime}}{\bf a}\cdot{\bf F}_{A_{n}^{\prime}}\Big{\}}$
$=$ $\frac{\tau}{m}\sum_{j\in a}{\bf a}\cdot{\bf f}_{j}.$
This cancels with the change in potential energy. Hence the superposed evolved momentum states maintain constant energy for the bosons originally in the momentum state ${\bf a}$. It is essential for this result that the force on the subsets $A$ and $A^{\prime}$ be reduced by the binomial coefficient.
The result for the evolution of the momentum of a subset of bosons, Eq. (2.8), says that Newton’s classical rate of change of momentum is shared amongst the ${}^{N_{\bf a}}C_{n_{A}}$ superposed states, which in total conserves energy. But since only one momentum configuration survives, we conclude that kinetic energy is lost to the environment via entanglement. This result explains physically the conditional transition probability derived next.
II.3 Adiabatic Stochastic Transition
@@
We now explicitly take the momenta to be quantized, with ${\bf p}$ being a $3N$-dimensional vector integer multiple of $\Delta_{p}$.
As in the preceding §II.2, we consider the transition of a subset $A$ of bosons in the momentum state ${\bf a}$ that moves rigidly according to the shared non-local force acting on the subset. The size distribution for the subsets will be derived in the following §II.4.
The present rigid subset model allows the momentum state to break up and the bosons ties to rearrange according to the state that they occupy after each time step or even after each successful transition. This shifts from the occupancy picture to the configuration picture, and therefore changes in occupation entropy must be taken into account. The configuration probability density is (Attard 2025)
$\wp({\bf\Gamma})=\frac{1}{Z}e^{-\beta{\cal K}({\bf p})}e^{-\beta U({\bf q})}\prod_{\bf a}N_{\bf a}!,$ (2.11)
where $\beta=1/k_{\mathrm{B}}T$ is the inverse temperature, ${\cal K}({\bf p})$ is the kinetic energy, $U({\bf q})$ is the potential energy, the commutation function has been neglected, and the occupancy of the momentum state ${\bf a}$ is $N_{\bf a}=\sum_{j=1}^{N}\delta_{{\bf p}_{j},{\bf a}}$. Also, a point in quantized phase space is ${\bf\Gamma}=\{{\bf q},{\bf p}\}$, and the conjugate point with momenta reversed is ${\bf\Gamma}^{\dagger}=\{{\bf q},-{\bf p}\}$.
Consider a specific subset $A\in{\bf a}$ of $n_{A}$ bosons. The shared non-local force per boson in the subset is ${\bf F}_{A}=n_{A}^{-1}\sum_{j\in A}{\bf f}_{j}$. We seek the conditional transition probability of this subset to the neighboring momentum state in the direction of the $\alpha$ component of the force on the subset, namely from ${\bf a}$ to ${\bf a}^{\prime}_{\alpha}={\bf a}+\mbox{sign}(\tau F_{A,\alpha})\Delta_{p}\widehat{\bf x}_{\alpha}$. Microscopic reversibility gives the ratio of conditional transition probabilities,
$\frac{\wp({\bf\Gamma}^{\prime}|{\bf\Gamma};\tau)}{\wp({\bf\Gamma}^{\dagger}|{\bf\Gamma}^{\prime{\dagger}};\tau)}$ $=$ $\frac{\wp({\bf\Gamma}^{\prime})}{\wp({\bf\Gamma})}$ (2.12)
@@
(This does not include the size probability distribution of the subsets; see §II.4.) By inspection, this is satisfied by a conditional transition probability with the general form
@@
$\Rightarrow$ $\lambda_{A,\alpha}\frac{n_{A}!(N_{{\bf a}}-n_{A})!}{N_{{\bf a}}!}$
@@
×$\times\left\{1-\frac{n_{A}\beta\Delta_{p}}{2m}\mbox{sign}(\tau{F}_{A,\alpha})a_{\alpha}+\frac{\beta\tau a_{\alpha}}{2m}\sum_{j\in A}{f}_{j,\alpha}\right\}.$
The exponential of the change in energy has been linearized here, although this is unnecessary and might in fact be ill-advised for larger subsets. In the particular solution that is the final equality, $r=1$ has been chosen. Likewise, $g(n_{A})=n_{A}!$ has been chosen (see §II.2; but similar values for the viscosity result from $g(n_{A})=\delta_{n_{A},1}$ (see §II.4), $g(n_{A})=1$, and the rigid state model $g(n_{A})=\delta_{n_{A},N_{\bf a}}$). This gives the known conditional transition probabilities for a rigid state transition, $n_{A}=N_{\bf a}$.
With this conditional transition probability, the average rate of change of momentum in the direction $\alpha$ for the subset $A\in{\bf a}$ to leading order is
@@
$\left\langle\dot{p}_{A}^{0}\right\rangle=\frac{n_{A}\Delta_{p}}{\tau}\mbox{sign}(\tau F_{A,\alpha})\lambda_{A,\alpha}\frac{n_{A}!(N_{{\bf a}}-n_{A})!}{N_{{\bf a}}!}.$ (2.14)
The classical regime is defined as $N_{{\bf a}}=n_{A}=1$, and in order to satisfy Newton’s second law of motion in this case we must have
$\lambda_{A,\alpha}\equiv\frac{|\tau F_{A,\alpha}|}{\Delta_{p}}.$ (2.15)
In the classical regime ${\bf F}_{\bf a}={\bf F}_{A}={\bf f}_{j}$ for $j\in A\in{\bf a}$.
With this result for $\lambda_{A,\alpha}$ and the conditional transition probability, Newton’s second law is not satisfied in the quantum condensed regime. (As a consequence neither energy nor momentum are conserved, which is not unexpected as these results apply to an open quantum subsystem; see §II.2.) Specifically, this says that for a highly occupied momentum state the rate of change of momentum for most subsets is much less than the classical prediction for a given applied force. The greatest reduction occurs when $n_{A}=N_{\bf a}/2$, in which case it is ${\cal O}(2^{-N_{\bf a}})$. Typically for low lying momentum states $N_{\bf a}={\cal O}(10^{2})$, and so this reduction is quite significant, exponentially small in fact. If the entire state attempts the transition as a rigid body with $n_{A}=N_{\bf a}$, then the combinatorial coefficient is unity and there is no reduction from the classical law. If an individual boson attempts the transition, $n_{A}=1$, then its rate of change of momentum is reduced by a factor of $N_{\bf a}^{-1}$ from the value given by Newton’s second law of motion.
In summary, for a given shared non-local force and a given subset, transitions from a momentum state are more frequent if the subset is the entire momentum state than if it consists of a fraction of the bosons in the state. But there are $(N_{\bf a}-1)(N_{\bf a}-1)!$ ways of choosing subsets that are not the entire state. Hence the vast majority of uniformly chosen subsets result in an exponentially reduced rate of change of momentum. But of course this all depends on the size distribution of the subsets, and this is derived in the following §II.4.
Since superfluid flow consists selectively of bosons in highly occupied states, this says that the rate of change of their momentum is exponentially reduced. It follows that there is a direct connection between the conservation of entropy on an adiabatic trajectory, the reduction in the rate of change of momentum in the condensed regime, and the loss of shear viscosity in superfluidity.
A second contribution to the reduction in the rate of change of momentum of a subset is the shared non-local force of the subset, $n_{A}{\bf F}_{A}=\sum_{j\in A}{\bf f}_{j}$. This shared non-local force is a consequence of the permutations of the bosons in the momentum state. The magnitude of the change due to this for the momentum state scales with $n_{A}^{1/2}$, whereas the dissipative effects on the classical shear viscosity of individual momentum changes scale with $n_{A}$. This effect is largest for the largest subsets, but is non-existent for individual transitions (see next). This non-local sharing effect reduces the rate of change of momentum in shear flow from its classical value. The combinatoral effect of the occupation entropy and the effect of the non-local sharing of the forces explain at the molecular level the reduction in viscosity in the condensed superfluid.
II.4 Distribution of Subset Size
Let $\wp(n_{A}|N_{\bf a})$ be the probability of a subset of size $n_{A}$ irrespective of its composition being involved in a transition from the momentum state containing $N_{\bf a}$ bosons. We now incorporate this into the unconditional transition probability. As in the preceding section, microscopic reversibility yields
$\wp({\bf\Gamma}^{\prime}|{\bf\Gamma};n_{A},\tau)\wp({\bf\Gamma})\wp(n_{A}|N_{\bf a})$ (2.16)
$=$ $\wp({\bf\Gamma}^{\dagger}|{\bf\Gamma}^{\prime{\dagger}};n_{A},\tau)\wp({\bf\Gamma}^{\prime{\dagger}})\wp(n_{A}|N_{{\bf a}_{\alpha}^{\prime}}^{\prime}).$
In the conjugate system, $N_{-{\bf a}}=N_{{\bf a}}$. We write $N_{\bf a}^{\prime}=N_{\bf a}-n_{A}$ and $N_{{\bf a}_{\alpha}^{\prime}}^{\prime}=N_{{\bf a}_{\alpha}^{\prime}}+n_{A}$.
The current conditional transition probability given in §II.3 remains valid if the size probability cancels here,
$\wp(n_{A}|N_{\bf a})=\wp(n_{A}|N_{{\bf a}_{\alpha}^{\prime}}^{\prime}).$ (2.17)
In general $N_{\bf a}\neq N_{{\bf a}_{\alpha}^{\prime}}^{\prime}$, and in view of the fact that the respective normalization constants in general depend upon each occupancy, and also the facts that $N_{\bf a}\geq 1$ and $N_{{\bf a}_{\alpha}^{\prime}}^{\prime}\geq 1$, the only way to ensure equality here is to choose
$\wp(n_{A}|N_{\bf a})=\delta_{n_{A},1}.$ (2.18)
This is equivalent to choosing $g(n_{A})=\delta_{n_{A},1}$ in Eq. (II.3), so that the conditional transition probability for boson $j\in{\bf a}$ becomes
$\wp_{j}({\bf a}^{\prime}_{j\alpha}|{\bf a},N_{\bf a})$